OpenStudy (anonymous):

What is the slope of the line 2x – 5y = 12 ? A. 2/5 B. 5/2 C. 2 D. -12/5

OpenStudy (uri):

Slope formula doe.

OpenStudy (uri):

y=mx+b

OpenStudy (anonymous):

@uri is b" the slope ?

OpenStudy (uri):

B the option or in the formula?

OpenStudy (anonymous):

the formula , but I found out .

OpenStudy (uri):

We have to add 5y both sides.

OpenStudy (uri):

2x-5y+5y=3+5y

OpenStudy (anonymous):

5y=-2x+12

OpenStudy (uri):

Now simplify it.

OpenStudy (anonymous):

divide everything by -5

OpenStudy (uri):

5y=2x-3.

OpenStudy (anonymous):